# Functional Brain-to-Brain Transformation with No Shared Data

### v1

@@ 5.2.1 Training brain-to-brain transformations with no shared data
The brain-to-brain transformation at the core of our method is implemented via a simple neural architecture featuring a single linear layer, optimized with an L2 regularization loss (detailed below). The image-to-fMRI encoder employed in our work is based on the framework outlined in [30]. It encompasses a pre-trained VGG network augmented with additional convolution layers and a non-linear layer to project data into the fMRI space. Prior to training the transformation network $T$, we performed separate training for the visual encoder of each subject individually (on their individual data), using the procedure recommended in [30] (refer to Figure 2.a). Notably, during transformation training, the encoder weights remained fixed.
@@
The transformation $T$ serves the purpose of mapping fMRI patterns from one subject (Subject1) to their corresponding patterns in another subject (Subject2). Leveraging the predictive capability of the pre-trained encoders for forecasting fMRI responses to images, we are capable of generating corresponding fMRI patterns for “non-shared” and “external” images. These are used to train the transformation $T$, without any shared data between the two subjects.
Our fMRI loss function adopts the mean square error loss between fMRIs:
$\mathcal{L}_{fMRI}(\hat{x},x)=||\hat{x}-x||_{2}$
where $x$ is the target fMRI, and $\hat{x}$ is the predicted fMRI after applying the transformation $T$.
@@
$\mathcal{L}_{S}=\mathcal{L}_{fMRI}({T}_{1,2}({F}_{1}),{F}_{2})$
where ${F}_{1}$ represents the recorded fMRI of Subject1, and ${F}_{2}$ represents the recorded fMRI of Subject2, while observing the same image. ${T}_{1,2}$ denotes the transformation model mapping from Subject1 to Subject2. We employ the fMRI loss of the transformation to compare the transformed fMRI with the actual measured fMRI.
@@
$\mathcal{L}_{NS}=\mathcal{L}_{fMRI}(T_{1,2}({F}_{1}),E_{2}({I}_{1}))\ \ or\ \ \mathcal{L}_{NS}=\mathcal{L}_{fMRI}(T_{1,2}(E_{1}({I}_{2})),{F}_{2})$
where ${F}_{1}$ and ${I}_{1}$ correspond to the recorded fMRI of Subject1 and the corresponding image (the image viewed exclusively by Subject1). We transform Subject1’s fMRI signals and compare them to the encoded fMRI signals derived from the image seen by Subject1(using Subject2’s encoder). In the same manners, ${F}_{2}$ and ${I}_{2}$ correspond to the recorded fMRI of Subject2 and the corresponding image (the image viewed exclusively by Subject2). We transform the encoded fMRI signals derived from the image seen by Subject2(using Subject1’s encoder) and compare them to Subject2’s fMRI signals $F_{2}$.
@@
It is worth noting that the number of shared and non-shared fMRI data is constrained by the available data, while the number of external images is essentially limitless. In occasions where we train the transformation without any shared data, either within or between datasets, we omit the first loss term $\mathcal{L}_{S}$ and employ only the other two loss components ($\mathcal{L}_{NS}$ and $\mathcal{L}_{Ext}$).
The coefficients governing the three loss terms and the regularization factor are established through a validation process. We separate the fMRI data into 3 sets: Train (80%), Validation(10%) and Test (10%). Subsequently, we select the optimal coefficients by assessing the transformation model performance on the validation set.
5.2.2 Improving Image-to-fMRI encoding of one subject using another subject’s fMRI
@@
Our comprehensive loss function comprises three components corresponding to different data types (shared, non-shared, and external). In each scenario, we can choose to use all or only some of these components together. Additionally, we include an encoding loss, which pertains to the student encoder and its available data. Lastly, a regularization term $\mathcal{L}_{Reg}$ employs L2 regularization on the weights of the linear transformation layer. Our loss function can be represented as follows:
@@
$\mathcal{L}_{Enc}=\mathcal{L}_{fMRI}(E_{Student}({I}_{Student}),{F}_{Student})$
@@
$\mathcal{L}_{NS}=\mathcal{L}_{fMRI}(E_{Student}({I}_{Teacher}),{F}_{Teacher})$
@@
$\mathcal{L}_{Ext}=\mathcal{L}_{fMRI}(E_{Student}({I}),E_{Teacher}({I}))$
The choice of the “teacher” subject is a crucial step in our method. We select the “teacher” subject based on its encoder’s performance when individually trained on its own data and evaluated on a validation set. The subject with the best-performing encoder is chosen for this role. Notably, the “teacher” subject can be selected from the same dataset or even from a different dataset, depending on the specific application.
To determine the optimal coefficients for the loss functions and regularization, we employ a validation procedure, as outlined in a previous section.
When applying our method to the “Vim1” dataset and the NSD dataset, we encountered a few challenges. These datasets differ in terms of color representation, with one being grey-scale and the other in RGB format. To bridge this domain gap, we converted the external and non-shared images to grey-scale, minimizing disparities between them.

### v2

@@ 5.4 Training brain-to-brain transformations with no shared data
The brain-to-brain transformation at the core of our method is implemented via a simple neural architecture featuring a single linear layer mapping all voxels of one subject onto each voxel of the other subject, optimized with an L2 regularization loss (detailed below). Prior to training the transformation network $T$, we performed separate training for the visual encoder of each subject individually (on their individual data), using the procedure recommended in [32] (refer to Figure 2.a). Notably, during transformation training, the encoder weights remained fixed.
@@
The transformation $T$ serves the purpose of mapping fMRI patterns from one subject (Subject1) to their corresponding patterns in another subject (Subject2). Leveraging the predictive capability of the pre-trained encoders for forecasting fMRI responses to images, we are capable of generating corresponding fMRI patterns for “non-shared” and “external” images. These are used to train the transformation $T$, without any shared data between the two subjects. Our fMRI loss function adopts the mean square error loss between fMRIs. The loss is calculated between the target measured fMRI of Subject2 and the predicted fMRI after applying the transformation $T$ from Subject1 to Subject2.
Losses.
@@
$\mathcal{L}_{S}=\mathcal{L}_{fMRI}({T}_{1,2}({\mathcal{F}}_{1}(\mathcal{I}_{Shared})),\mathcal{F}_{2}(\mathcal{I}_{Shared}))$
where $\mathcal{F}_{1}$ represents the recorded fMRI of Subject1, and $\mathcal{F}_{2}$ represents the recorded fMRI of Subject2, while observing the same image $\mathcal{I}_{Shared}$. ${T}_{1,2}$ denotes the transformation model mapping from Subject1 to Subject2. We employ the fMRI loss of the transformation to compare the transformed fMRI with the actual measured fMRI.
@@
$\mathcal{L}_{NS}=\mathcal{L}_{fMRI}(T_{1,2}(\mathcal{F}_{1}(\mathcal{I}_{N_{s_{1}}})),E_{2}(\mathcal{I}_{N_{s_{1}}}))\ \ or\ \ \mathcal{L}_{NS}=\mathcal{L}_{fMRI}(T_{1,2}(E_{1}(\mathcal{I}_{N_{s_{2}}})),\mathcal{F}_{2}(\mathcal{I}_{N_{s_{2}}}))$
where $\mathcal{F}_{1}(\mathcal{I}_{N_{s_{1}}})$ and $\mathcal{I}_{N_{s_{1}}}$ are the measured fMRI of Subject1 and the corresponding image (the image viewed exclusively by Subject1). We transform Subject1’s fMRI signals and compare them to the encoded fMRI signals derived from the image seen by Subject1 (using Subject2’s encoder). In the same manners, $\mathcal{F}_{2}(\mathcal{I}_{N_{s_{2}}})$ and $\mathcal{I}_{N_{s_{2}}}$ correspond to the recorded fMRI of Subject2 and the corresponding image (the image viewed exclusively by Subject2). We transform the encoded fMRI signals derived from the image seen by Subject2 (using Subject1’s encoder) and compare them to Subject2’s fMRI signals $\mathcal{F}_{2}$.
@@
It is worth noting that the number of shared and non-shared fMRI data is constrained by the available data, while the number of external data is essentially limitless. In occasions where we train the transformation without any shared data, either within or between datasets, we omit the first loss term $\mathcal{L}_{S}$ and employ only the other two loss components ($\mathcal{L}_{NS}$ and $\mathcal{L}_{Ext}$).
Training Technical Details.
The training is conducted using scikit-learn’s Ridge regression model using the different kind of data types. The samples are weighted according to the loss coefficients mentioned above ($\alpha_{S}$, $\alpha_{NS}$ and $\alpha_{Ext}$). We separate the fMRI data into 3 sets: Train (80%), Validation (10%) and Test (10%). Subsequently, we select the optimal regularization term by assessing the transformation model performance on the validation set. When training with non-shared fMRI data we also used 6000 external images. We used $\alpha_{S}=1,\alpha_{NS}=0.5$ and $\alpha_{Ext}=0.1$.
5.5 Improving Image-to-fMRI Encoding via B2B Transformation
@@
Losses.
Our comprehensive loss function comprises three components corresponding to different data types (shared, non-shared, and external) used for training both the student encoder and the transformation. In each scenario, we can choose to use all or only some of these components together. Additionally, we include an encoding loss, which pertains to the student encoder and its available data. Lastly, a regularization term $\mathcal{L}_{Reg}$ employs L2 regularization on the weights of the linear transformation layer. Our loss function can be represented as follows:
@@
$\mathcal{L}_{Enc}=\mathcal{L}_{fMRI}(E_{Student}({I}_{Student}),{F}_{Student}({I}_{Student}))$
@@
$\mathcal{L}_{NS}=\mathcal{L}_{fMRI}(E_{Student}({I}_{Teacher}),T_{T,S}({F}_{Teacher}({I}_{Teacher})))$
@@
$\mathcal{L}_{Ext}=\mathcal{L}_{fMRI}(E_{Student}({I}_{Ext}),{T}_{T,S}(E_{Teacher}({I}_{Ext})))$
Training Technical Details.
The training is conducted using the Adam optimizer for 30 epochs with an initial learning rate of 5e-4 (employing step decay) to update the weights of both the encoder and the transformation model. Each training batch consists of 32 examples from each type: 32 image-fMRI pairs from the student subject, 32 image-fMRI pairs from the teacher subject for non-shared data, 32 pairs for shared data (if available) and another 32 external images (which are encoded using both the teacher and student encoders). In each batch, different pairs from the teacher and different external images are sampled to ensure variety and comprehensive training. To determine the optimal coefficients for the loss functions and regularization, we employ a validation procedure, as outlined in a previous section. When applying our method to the “Vim1” dataset and the NSD dataset, we encountered a few challenges. These datasets differ in terms of color representation, with one being grey-scale and the other in RGB format. To bridge this domain gap, we converted the external and non-shared images to grey-scale, minimizing disparities between them.
The selection of the teacher subject is a critical step in our method. We choose the teacher based on its effectiveness in enhancing the encoding performance of other subjects on a validation set. Importantly, the teacher can be chosen from within the same dataset or from a different dataset, depending on the specific requirements of the application.
5.6 Decoding model
In this study, we employed a decoding process to reconstruct images from fMRI data following the completion of the transformation. Our decoding process combines two distinct methods, which are detailed in [32] and [39]. First, we utilized the model introduced in [32]. This model, when provided with an fMRI input, generates a reconstructed image that closely aligns with the structural aspects of the original image. However, the output may not appear entirely natural. To address this limitation, we incorporated a novel decoding model from [39]. This model leverages a diffusion-based approach to produce highly realistic natural images from fMRI data. Nevertheless, it may not preserve the fine structural details as effectively. To strike a balance between fidelity to the original image and realism, we adopted a two-step decoding strategy. Initially, we employed the first model to generate a reconstructed image, which faithfully captures the structural aspects of the original but might lack naturalness. Then, we used this initial reconstruction as the starting point for the second model, which refines the image to make it both faithful to the original and visually realistic. This combined approach allows us to achieve a compelling balance between structural accuracy and natural appearance in the reconstructed images.
5.7 Statistical tests
The statistical test that was used throughout this work is a two-way paired permutation test with a one-sided alternative. The number of permutations that was used is 100000, and the statistic calculated is the mean of the differences. This was done using the stats.permutation_test function in the python package scipy. In figures, * marks $p<0.05$, ** marks $p<0.01$ and *** marks $p<0.001$, and the dotted lines mark which models were compared.
@@ 6 Data availability
7 Code availability
The code for both the brain-to-brain transformation and the enhanced encoder methods, along with scripts used to create the figures for the paper, is available on GitHub at
https://github.com/navvewas/Brain-Transformation-with-No-Shared-Data.git
8 Declaration of generative AI and AI-assisted technologies in the writing process
During the preparation of this work the authors used ChatGPT in order to improve the readability and language of the manuscript. After using this tool/service, the authors reviewed and edited the content as needed and take(s) full responsibility for the content of the published article.
9 Acknowledgments
This research was supported by the European Research Council (ERC) under the Horizon program, grant number 101142115.
Supplementary Material
S1 Analyzing Encoder Enhancement
In this section, we analyze various factors that contribute to improving Image-to-fMRI encoder performance and offer deeper insights into the effectiveness of our method. Figure S1 consolidates four plots that dissect different aspects of our approach.
We first show that a real fMRI scan, when B2B-transformed from a source-subject to a target-subject, exceeds the quality of Image-to-fMRI encoding of the target-subject (and is particularly powerful if the target-subject has only a small number of examples to train its encoder). This assumption is validated in Figure S1.a, where the source subject was Subject4 of the GOD datasdet, and the 4 other GOD subjects (Subjecti, i=1,2,3,5) served as target subjects. The Image-to-fMRI encoder of the 4 target subjects were trained using 300 (25%) of their training examples. The figure shows that a real fMRI of Subject4, when B2B-transformed to a target subject, outperforms the emphencoded-fMRI obtained from the target-subject’s Baseline-encoder. The grey bars represent the performance of the Baseline-encoder, while the red bars show the mean Pearson correlation between the real fMRI of GOD Subject4 after being B2B-transformed into the space of the target-subject. The correlations are measured with respect to the real fMRI scan of the target-subject (evaluated on the GOD test-data). This indicates that the B2B-transformed fMRI can indeed be used to boost encoder performance of another subject (especially when that subject has few training data).
The quality of the “teacher” data is another critical factor in improving encoder performance. In Figure S1.b, we demonstrate that a subject with higher-quality data serves as a more effective teacher. For this analysis, we still used GOD subjects as students, but this time with NSD subjects as teachers (we used 7 NSD subjects as teachers, excluding one outlier with a significant gap between data quality and encoding performance). We trained each GOD subject with one (out of 7) NSD subject as a teacher on 300 examples. For each GOD subject, we obtained seven encoding correlations, corresponding to the seven NSD teachers. We perform this experiment for each of the 3 repeats and average them. The figure shows the mean and SEM for GOD Subject 1 (see Figure S12 for the remaining subjects). A clear trend emerges, with a correlation between teacher SNR and encoding performance.
In Figure S1.c, we first analyze the benefits of using a 7-Tesla NSD teacher versus a 3-Tesla GOD teacher, with the same number of teacher training examples (purple versus orange bars). The results show the performance of each GOD subject trained with 300 examples, enhanced via a teacher subject. We compare using Subject4 from GOD (purple) and Subject1 from NSD (red) as teachers, ensuring a fair comparison by using teacher data with the same number of available examples (1200 training pairs and three repetitions per example). The plot reveals that using NSD as the teacher consistently outperforms the GOD teacher across all four GOD student subjects. Additionally, when a large number (7100) of the NSD teacher training examples are used (orange), there is a noticeable performance boost, highlighting the importance of both high-quality teacher fMRI data and a larger number of examples in improving encoder performance.
Finally, we evaluate the impact of increasing the number of repetitions in the GOD dataset. The updated GOD dataset version includes five repetitions per example. As shown in Figure S1.d and in Figure S9, more repetitions lead to improved encoder performance. This is expected, as averaging more repetitions reduces noise in the fMRI, thus improving its signal-to-noise ratio (SNR). Interestingly, even without additional repetitions, our B2B method for enhancing encoders produces high performance results, suggesting that our approach could serve as an alternative to acquiring more data or repetitions.
Figure S1: Encoder Enhancement Analyzing: (a) The figure demonstrates that a real fMRI scan from Subject 4, after being B2B-transformed into the space of target subjects (Subjects 1, 2, 3, and 5), outperforms the Image-to-fMRI encoding of the target subjects’ Baseline encoders trained on 300 examples. Grey bars represent the Baseline-encoder performance, while red bars show the Pearson correlation between the B2B-transformed fMRI and the target subjects’ real fMRI scans, evaluated on GOD test data. This highlights the potential of B2B-transformed fMRI to enhance encoder performance, particularly for subjects with limited training data. (b) Teacher quality analysis: the correlation between teacher voxels mean SNR and encoding performance, shown for GOD Subject 1 using NSD subjects as teachers (mean and SEM across three repetitions). Higher-quality teacher data leads to better encoding performance showed by the postive correlation between the teacher SNR with the average mean correlations over repetitions ($r=0.85$, $R^{2}=0.735$, $p=0.006$). (c) Comparison of GOD student encoder performance with 1200 training examples enhanced by GOD or NSD teachers. Models trained with NSD teachers consistently outperform those trained with GOD teachers, with further improvement observed when all NSD examples are utilized. (d) Impact of repetitions: using more repetitions per example in the GOD dataset improves encoding performance, as shown for GOD subjects. Even without additional repetitions, our method achieves notable performance gains.
Subject Baseline Encoder + Shuffled Data + Random Data + Duplicated Data
1 0.1830 0.1416 0.1283 0.1725
2 0.2152 0.1701 0.2047 0.2153
3 0.1590 0.1131 0.1280 0.1626
4 0.2515 0.1947 0.2033 0.2478
5 0.0779 0.0638 0.0670 0.0813
Average 0.1773 0.1289 0.1462 0.1759
Table S1: Impact of Augmented Data on Encoder Performance: This experiment evaluates whether the improvement seen in the encoder’s performance, as demonstrated in the paper, is due to the addition of meaningful data from a teacher subject or simply the result of having more training data. To explore this, we present results for three different augmentation methods used to add 300 additional examples to the baseline model, which was initially trained with 300 original examples. These methods include: (i) shuffling voxel values within the original fMRI data to generate synthetic examples, (ii) creating random fMRI data with the same distribution as the measured data, and (iii) duplicating the original 300 examples. The results show that these simple augmentation methods fail to significantly improve performance, underscoring the importance of meaningful teacher-derived data for boosting encoder performance.
Figure S2: Transformation between NSD and GOD datasets: This figure presents two heatmaps showing the mean Pearson correlation for all subject pairs: one for transformations from GOD to NSD and the other from NSD to GOD. As there is no shared data between the datasets, we generated corresponding fMRIs for the target dataset’s test set using the encoder trained on the source dataset. These encoded fMRIs were then transformed to match the target subject and evaluated by comparing them to the real measured test fMRIs.
Figure S3: Pairwise Correlation and Retrieval for Transformations: The figure presents heatmaps for all NSD subject pairs, showing both correlation and retrieval accuracy for four different transformation methods: (i) anatomical mappings, (ii) transformations using 200 shared examples, (iii) transformations using 700 shared examples, and (iv) our method, which leverages all available non-shared data (6400 examples) and external images.
(a) Transformation of Subject1 to Subject2
(b) Transformation of Subject2 to Subject1
(c) Transformation of Subject3 to Subject4
(d) Transformation of Subject4 to Subject3
Figure S4: Evaluating brain-to-brain transformation with no shared data: In this figure, we conduct a quantitative assessment by calculating the mean Pearson’s r correlation for different transformations within the NSD dataset across various transformation models. We compare the resulting transformations when trained only with non-shared data (solid red line) to three other transformations: anatomical mapping to a common brain space (cyan dotted line), transformation trained with only 200 shared examples (dotted blue line), and transformation trained with 700 shared examples (dotted grey line). The remaining 300 “shared data” serve as “test data” for assessing and comparing the quality of all the learned transformations. The x-axis of the graph represents the number of non-shared examples available for training, where our transformation model is the only one that can use non-shared examples. (a): A single example of the Pearson’s r results for the transformation between Subject1 and Subject2 in the NSD dataset. (b): Same as ’a’, between Subject2 and Subject1. (c): Same as ’a’, between Subject3 and Subject4. (d): Same as ’a’, between Subject4 and Subject3.
(a) Mean of all transformations
(b) Transformation Subject1 to Subject2
(c) Transformation Subject2 to Subject1
(d) Transformation Subject3 to Subject4
Figure S5: Using non shared data alongside shared data to improve transformations: Across all figures, we make a comparison between a transformation trained with only 200 shared examples (shown as a dotted green line) and a transformation trained using our method, which includes non-shared examples in addition to the 200 shared examples (shown as a green line). The x-axis on each figure represents the quantity of non-shared examples utilized for training. (a): Percentage improvement in Pearson’s r for models using varying amounts of non-shared data (green line) compared to the transformation trained with 200 shared examples (dotted green line). The error bars represent the SEM over the pairs. (b): A single example of the Pearson’s r results for the transformation between Subject1 and Subject2 in the NSD dataset. (c): Same as ’b’, between Subject2 and Subject1. (d): Same as ’b’, between Subject3 and Subject4.
(a) Mean of all transformations
(b) Transformation Subject1 to Subject2
(c) Transformation Subject2 to Subject1
(d) Transformation Subject3 to Subject4
Figure S6: Using non shared data alongside shared data to improve transformations: Across all figures, we make a comparison between a transformation trained with only 700 shared examples (shown as a dotted green line) and a transformation trained using our method, which includes non-shared examples in addition to the 700 shared examples (shown as a green line). The x-axis on each figure represents the quantity of non-shared examples utilized for training. (a): Percentage improvement in Pearson’s r for models using varying amounts of non-shared data (green line) compared to the transformation trained with 700 shared examples (dotted green line). The error bars represent the SEM over the pairs. (b): A single example of the Pearson’s r results for the transformation between Subject1 and Subject2 in the NSD dataset. (c): Same as ’b’, between Subject2 and Subject1. (d): Same as ’b’, between Subject3 and Subject4.
Figure S7: Image Reconstruction from Transformed fMRI of Subject1 to Subject2: Subject1’s fMRI data was transformed to align with the fMRI space of Subject2. This transformed data was then decoded into an image using a pre-trained decoder specific to Subject2. The left column displays the original image presented to Subject1, while the second column illustrates the decoded images following anatomical transformation. The subsequent column presents the decoded image after employing our method, trained exclusively with non-shared examples.
Figure S8: Image Reconstruction from Transformed fMRI of Subject2 to Subject1: Subject2’s fMRI data was transformed to align with the fMRI space of Subject1. This transformed data was then decoded into an image using a pre-trained decoder specific to Subject1. The left column displays the original image presented to Subject2, while the second column illustrates the decoded images following anatomical transformation. The subsequent column presents the decoded image after employing our method, trained exclusively with non-shared examples.
Figure S9: Improving encoding of a subject from GOD 3-Tesla dataset using another high-quality subject: This figures compares the quality of the “student” baseline encoder (which is trained solely on the student’s own fMRI data), to the student encoder obtained when incorporating also data from the “teacher” during training. In each plot, the grey line represents the student baseline encoder model, the purple line symbolizes the encoder improvement obtained when using a “teacher” from the same dataset (e.g., Subject4 is the highest-quality subject in the GOD dataset). The orange line corresponds to the encoder enhancement achieved by utilizing a superior subject from another dataset (NSD), which has more scanned examples and higher fMRI resolution (7-Tesla machine). The quality of the resulting encoders is assessed through the mean Pearson correlation of all predicted fMRI voxels. The “number of training examples” refers to the number of the student’s own image-fMRI pairs of examples used for training the “student” encoder, whereas the “teacher” subject (drawn from either GOD or NSD dataset) has been trained using all its own available examples and remains unchanged during the training of the student encoder. All plots present specific subject encoding results quantified using the mean Pearson correlation of all voxels across different models, based on the number of training examples.
Figure S10: Improving classification accuracy (of never-before-seen image classes) of a 3-T GOD subject, using its enhanced encoder via other subject: The classification process, involves using an encoder to convert images without any fMRI measurements into fMRI patterns, even for semantic classes that the subject has never seen before. For each class, we feed 100 different images into the encoder, generating 100 corresponding fMRI patterns. These patterns are then averaged for each class, creating an “average fMRI class representation” for each category. Consequently, when given a new test fMRI pattern, we can classify it by comparing it to these averaged fMRI class representations, to determine the correct class.
Figure S11: Image Classification from fMRI of 3-T GOD subjects: This plot presents the classification results employing this approach, considering different modes of encoder training. We compare the baseline encoder (trained only on its subject-specific fMRI data) with encoders improved by another subject (“teacher”), either from the GOD dataset or using the NSD 7-Tesla dataset. The results presented are the mean accuracy of all subjects with the SEM error over different subjects and individual subject result with mean and SEM over 3 training repetitions.
Figure S12: Teacher quality analysis: the correlation between teacher SNR and encoding performance, shown for GOD Subjects using NSD subjects as teachers (mean and SEM across three repetitions). Higher-quality teacher data leads to better encoding improvements.
(Shared, Non-Shared) Subject
1 2 3 4 5 6 7 8
(100, 0) 0.2124 0.2484 0.1424 0.1513 0.2274 0.1449 0.1130 0.1052
(200, 0) 0.2745 0.3094 0.1766 0.1954 0.2827 0.1882 0.1532 0.1131
(400, 0) 0.3204 0.3585 0.2226 0.2302 0.3213 0.2207 0.2018 0.1368
(700, 0) 0.3608 0.3902 0.2542 02514 0.3651 0.2487 0.2362 0.1625
(700, 900) 0.4272 0.4591 0.3172 0.3272 0.4334 0.3278 0.3042 0.2197
(700, 2500) 0.4581 0.4915 0.3615 0.3672 0.4681 0.3685 0.3513 0.2529
(700, 5700) 0.4753 0.5033 0.3819 0.3818 0.4837 0.3880 0.3651 0.2775
(700, 6400) 0.4752 0.5040 0.3761 0.3862 0.4796 0.3862 0.3677 0.2808
Table S2: NSD encoders correlation: Mean correlation values for each subject and example number pair (shared, non-shared) for all subjects in the NSD dataset.